# Learning Physics From Video: Unsupervised Physical Parameter Estimation for Continuous Dynamical Systems

### v1

@@ Abstract
Extracting physical dynamical system parameters from videos is of great interest to applications in natural science and technology. The state-of-the-art in automatic parameter estimation from video is addressed by training supervised deep networks on large datasets. Such datasets require labels, which are difficult to acquire. While some unsupervised techniques –which depend on frame prediction– exist, they suffer from long training times, instability under different initializations, and are limited to hand-picked motion problems. In this work, we propose a method to estimate the physical parameters of any known, continuous governing equation from single videos; our solution is suitable for different dynamical systems beyond motion and is robust to initialization compared to previous approaches. Moreover, we remove the need for frame prediction by implementing a KL-divergence-based loss function in the latent space, which avoids convergence to trivial solutions and reduces model size and compute.
@@ 1 Introduction
Figure 1: We propose a novel unsupervised approach to physical parameter estimation from videos. The black squared images are video frames with different states of a white pendulum. Starting from a frame at time $t$ (center) parameter estimation techniques typically employ an encoder, which estimates the dynamical states $z_{t}$ at time $t$. A physics block (Physics ODE), with learnable parameters, then solves the governing equations of the dynamical system to predict future dynamical states $\hat{z}_{t+1}$, which takes place in the latent representation (blue lines). Previous state-of-the-art methods (left) then design decoders to reconstruct the next frame at time $t+1$ and use reconstruction loss ($\mathcal{L}$, purple left) to train the physics ODE block. In contrast, our method (right) completely avoids the need for a decoder by leveraging a loss function in the latent space ($\mathcal{L}$, right). Our loss (blue) function minimizes the distance between the estimated states $\hat{z}_{t+1}$ and $z_{t+1}$.
@@
Fitting governing equations is an inverse problem [2], which often requires either using additional sensors to directly measure system states. Instead, doing measurements from a video allows to avoids additional sensors, yet, requires manually labelling pixels or video frames which is time-consuming and expensive. Therefore, automated and unsupervised methods are needed to extract dynamics from videos and accurately estimate physical parameters [6, 15, 18, 17, 32].
@@
Recent work addressed parameter estimation from video by deep learning [6, 7, 37] or reinforcement learning [3]. For instance, proposed supervised learning methods rely on datasets with extensive and high precision labels which are exceedingly difficult to obtain [1, 4, 26, 28, 35]. To avoid labeling, current unsupervised methods for estimating physical parameters build on encoder-decoder network designs: reconstructing video frames from low-dimensional representations. The frame reconstruction is a mere by-product of the paramter estimation, and leads to overly complex solutions which are difficult to train [15, 18]. Along with this, current solutions [15, 18, 20, 38, 43] are constrained to motion dynamics, excluding a wide variety of systems like dynamics related to brightness, colour, and deformations, among others [15, 18].
Our work proposes an unsupervised learning method to solve the inverse problem using videos of a dynamical system with known, continuous governing physics equations. Our method can be implemented for different dynamical systems beyond motion problems. Thus, unlike previous approaches, we present an evaluation of the latent space of our model in multiple dynamical systems. We show that our unsupervised model successfully fits the dynamics and effectively generalizes to unseen future frames. In addition, we bypass frame reconstruction by calculating the loss simply in the latent space, eliminating the need for a decoder. Our approach is thus fast, less resource-intensive and more robust to initial conditions compared to existing methods.
Our approach is visualised in Figure 1, and our main contributions are summarized as:
@@
  • Allowing inverse problems beyond motion, such as appearance and scaling.
  • An unsupervised latent space loss to avoid expensive video frame reconstruction;
  • Easy model training with robustness against different initial conditions.
Our code and datasets will be made available on GitHub.
@@ 2 Related Work
The relationship between physics and deep learning is symbiotic. Physics inspired segmentation [25] and generative models [34, 33] as well as the design of new architectures [12]. Likewise, deep learning is used to study, understand and create new physics from data [6, 7, 17, 37]. Techniques like physics-informed neural networks (PINN) [21, 31] or Lagrangian neural networks (LLN) [8, 27] are designed to solve inverse problems. Yet, PINNs are usually constrained to initial conditions, boundary conditions, and time reference [21, 11, 29]. Moreover, these methods are supervised and require labeled data. Because obtaining labeled data in inverse problems is expensive or even infeasible [1] we avoid the need for labels and offer an unsupervised method, following [1, 4, 26, 28, 35]. Our work focuses on the the inverse problem of physical parameter estimation from video using unsupervised machine learning.
Learning Physics from Video
@@
Work on learning physics from recorded videos focus on frame prediction [5, 10, 13, 24] and do not concentrate on estimating physical parameters from video. Existing research on extracting physical information from video [40] estimates parameters such as the friction coefficient, or the value of the dynamical state variable such as the position or velocity. These methods are supervised and require labelled datasets with access to the dynamical variables or parameters’ ground truth [40, 9, 27, 31, 39, 37, 42, 41]. Moreover, these methods mainly focus on motion problems and are based on systems similar to interaction networks [5, 37, 36]. Some methods [38, 19] aim to parameterize and solve differential equations to simulate the deformations. However, dynamical systems are not limited to motion, and we here propose a method that goes beyond motion which we illustrate with use-cases for appearance changes, and scaling.
Unsupervised Parameter Estimation from Video
@@
Existing unsupervised work use frame representation and prediction with physics priors of known governing equations with unknown parameters to estimate [38, 18, 15, 43, 20]. Some approaches [20, 43] use variational auto-encoders (VAE)[23] with a physics engine between the encoder and the decoder to estimate parameters; however, the reconstructions are poor, constraining the model to simple motion problems. The most similar work to ours are [15, 18] where they estimate parameters from a single video without annotations, and we experimentally compare against them. Our method is favourably compared in terms of robustness to initializations; we examine the latent space and extension to different dynamics.
Baselines
@@
As a strong, unsupervised parameter estimator, Jaques et al. [18] uses a traditional auto-encoder with a physics engine in the latent space to reconstruct inputs and generate future frame predictions. The model also incorporates a U-Net in the encoder to learn segmentation masks for the object of interest in an unsupervised way; the capability to learn this mask is linked to the spatial transformer (ST) they use in the decoder similar to [16]. During inference, the spatial transformer is used to perform affine transformations on the mask to ’move’ the object in future frame predictions. This limits the application of [18] to affine motion dynamics.
@@
On the other hand, Hofherr et al. [15] uses a differentiable ODE solver to estimate the parameters. This model also uses a spatial transformer, but at the pixel level, the object pixels are displaced using the prediction made by the ODE solver. The model needs to be trained, along with the use of masks, to learn which pixels should be translated. Finally, using frame reconstruction to achieve parameter estimation is challenging and makes the network slow to train since reconstructing frames from low dimensional data (i.e., a set of positions and velocities) is an ill-defined problem. Therefore, [15, 18, 38] had to limit their scope to using a mask and a spatial transformer [16], excluding dynamical systems with changes in intensity and colour, deformations and non-uniform scaling among others which we explicitly allow in our paper.
3 Methods
@@
Figure 2: Method overview. A video recording of an object with a periodic brightness change (bottom) displays dynamics $z_{\text{real}}$ with sampling period $\delta t$. Each frame is mapped by the encoder $E_{\theta}(\cdot)$ to the unsupervised latent representation $z_{t}$. The physics block $P_{\gamma}(\cdot)$ generates a prediction of the future step $\hat{z}_{t+1}$, which we compare to the encoded representation $z_{t+1}$ of the corresponding frame. Top-right: Loss function of our model; the first term ensures the prediction fits with the encoding, while the second expression controls the variance of $z$.
Our approach aims to estimate the parameters of a known governing equation from a video recording; since it is unsupervised, the training set consists of unannotated frames with a known frame rate denoted as $\delta t$. While existing methods which depend on frame prediction use complete auto-encoder architectures, our method is composed only of a simple encoder and a physics block. Figure 2 shows our approach altogether: first, the relation between the system dynamics and the frames; second, how the encoder processes the video frame by frame to obtain the latent space representations; and finally, how our physics block makes predictions to be used in our unsupervised loss function.
Scope
@@
We study systems represented by autonomous differential equations since those models just depend of the state variable which is captured in video, and it is not affected by external forces and takes the form:
$z^{(n)}+\gamma_{n-1}z^{(n-1)}+...+\gamma_{1}z^{(1)}+\gamma_{0}z=0,$ (1)
@@
which is an $n^{\textrm{th}}$-order system, where $z$ is the time-dependent state variable or "dynamic variable", $z^{(k)}$ with $k=1,2,\dots n$ is the $k^{\textrm{th}}$-derivative of $z$ with respect to time $t$ and $\gamma_{i}$ with $i=0,1,..n-1$ are the parameters of the equation we want to estimate.
@@
In the following Proof-of-concept we first consider a second-order differential equation:
@@
Encoder: The encoder is a network $E_{\theta}(x)$ that maps images $x\in\mathbb{R}^{w\times h\times c}$ to the state variable $z\in\mathbb{R}^{d}$, where $d$ is the number dynamic variables to study. We used an MLP with three layers and ReLU as an activation function for the mapping.
$E:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{d}$
$z_{t}=E_{\theta}(x_{t}).$ (3)
Physics block: Our physics block numerically solves the differential equation using Euler’s method:
@@
$\hat{z}_{t+1}=z_{t}+\delta t\left(z^{(1)}_{t}-\delta t(\gamma_{1}z^{(1)}_{t}+\gamma_{0}z_{t})\right)$ (7)
$P:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
$\hat{z}_{t+1}=P_{\gamma}(z_{t},..z_{t-n};\bf{\gamma})$ (8)
where $\gamma_{i}$ are learnable parameters and the predicted latent space for time $t+1$ is a function $P_{\gamma}(\cdot)$ of the latent representations of the $n$ previous frames. We use the notation $\mathbf{\hat{z}}$ for the latent space predicted using the function $P$, which is different from $\mathbf{z}$ predicted by the encoder.
Predictions: The encoder maps images $x_{t}\in\mathbb{R}^{w\times h\times c}$ to the state variable $z_{t}\in\mathbb{R}^{d}$ for all time steps $t\in[0,T]$ leading to $\hat{z}$ with dimensionality $\mathbb{R}^{T\times d}$ for all input frames. In particular, we have the following two vectors:
@@
Loss function: The first goal of the loss function is simply to minimize the difference between the predictions $\mathbf{z}$ and $\mathbf{\hat{z}}$ over a batch of size $M$. Therefore, our loss function is given first by:
$\mathcal{L}_{1}=\frac{1}{M}\sum_{i=1}^{M}(z_{i}-\hat{z}_{i})^{2}=\\ \frac{1}{M}\sum_{i=1}^{M}\Bigl{[}E_{\theta}(x_{i})-P_{\gamma}\bigl{(}E_{\theta}(x_{i-1}),\dots,E_{\theta}(x_{i-n})\bigr{)}\Bigr{]}^{2}$ (11)
However, one problem with this approach is the convergence to the trivial solution such that ${E_{\theta}(x)=0\;\forall x}$ and $P_{\gamma}(z)=0\;\forall z$. To avoid this problem, we want to induce variance in the encoder’s output: we propose a variational latent space and assume $z_{k}\in\mathcal{N}(\mu,\sigma^{2})$. The values of $\mu$ and $\sigma^{2}$ effectively define the range of $z$ since this consists of a re-normalization of the metric. Based on the VAE approach [23], we define the second part of the loss function using the Kullback-Leibler divergence (KL-divergence). Thus, $z_{k}$ is an element of the random variable $Z\sim\mathcal{N}(\mu_{z},\sigma_{z}^{2})$ and we want it to follow a particular distribution $Q\sim\mathcal{N}(0,1)$, then KL-divergence is given by:
$\mathcal{L}_{2}=\text{KL}(Z||Q)=-\frac{1}{d}\sum_{i=1}^{d}\bigl{(}1+\ln(\sigma_{z}^{2})-\mu_{z}^{2}-\sigma_{z}^{2}\bigr{)}$ (12)
@@
$\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}=\frac{1}{M}\sum_{i=1}^{M}(z_{i}-\hat{z}_{i})^{2}-\frac{1}{d}\sum_{i=1}^{d}\bigl{(}1+\ln(\sigma_{z}^{2})-\mu_{z}^{2}-\sigma_{z}^{2}\bigr{)}$ (13)
@@
Our loss function (Eq. 13) serves several purposes: First, $E_{\theta}(\cdot)$ maps images to random variables $z$ in latent space, and $L_{2}$ guarantees that $z$ is normally distributed. Then, the function $P_{\gamma}(\cdot)$ preserves the sense of order and relation between latent spaces from different frames. Since we are analyzing a single video, $z$ contains information about the temporal properties of the sequence of frames. Finally, $L_{1}$ makes the model consistent since the physical predictions made by the physics block $P_{\gamma}(\cdot)$ are generated by $E_{\theta}(\cdot)$.
@@
Training: For parameter estimation in the physics block $P_{\gamma}(\cdot)$, we used a learning rate proportional to the initial value $\gamma_{k}^{0}$ of the learnable parameter $\gamma_{k}$, where $\text{lr}_{(\gamma_{k})}\sim 10^{[\log_{10}|(\gamma_{k}^{0})|]}$. This approach provides sufficiently large step sizes at the beginning of training to escape of local minima. Further training details are in Appendix A.2.

### v2

@@ Abstract
Extracting physical dynamical system parameters from recorded observations is key in natural science. Current methods for automatic parameter estimation from video train supervised deep networks on large datasets. Such datasets require labels, which are difficult to acquire. While some unsupervised techniques–which depend on frame prediction–exist, they suffer from long training times, initialization instabilities, only consider motion-based dynamical systems, and are evaluated mainly on synthetic data. In this work, we propose an unsupervised method to estimate the physical parameters of known, continuous governing equations from single videos suitable for different dynamical systems beyond motion and robust to initialization. Moreover, we remove the need for frame prediction by implementing a KL-divergence-based loss function in the latent space, which avoids convergence to trivial solutions and reduces model size and compute. We first evaluate our model on synthetic data, as commonly done. After which, we take the field closer to reality by recording Delfys75: our own real-world dataset of 75 videos for five different types of dynamical systems to evaluate our method and others. Our method compares favorably to others. Code and data are available online: https://github.com/Alejandro-neuro/Learning_physics_from_video.
@@ 1 Introduction
Figure 1: We propose a novel unsupervised approach to physical parameter estimation from videos. Black squares are video frames with different states of a white pendulum. Starting from a frame at time $t$ (center) an encoder estimates the dynamical states $z_{t}$. A learnable physics block (Physics ODE) solves the dynamical system equations to predict future states $\hat{z}_{t+1}$ in latent space (blue lines). Previous state-of-the-art methods (left) then use decoders and a reconstruction loss ($\mathcal{L}$, purple left) to train the physics ODE block. In contrast, our method (right) completely avoids the need for a decoder by leveraging a loss function in the latent space ($\mathcal{L}$, blue right). Our loss function minimizes the distance between the estimated states $\hat{z}_{t+1}$ and $z_{t+1}$.
@@
Fitting governing equations is an inverse problem [2], which often requires using additional sensors to directly measure system states. Video-based measurements can eliminate the need for additional sensors, yet, require manually labelling pixels or video frames which is time-consuming and expensive. Therefore, automated and unsupervised methods are key to extract dynamics from videos and accurately estimate physical parameters [7, 17, 20, 19, 38].
@@
Recent work addressed parameter estimation from video by deep learning [7, 8, 44] or reinforcement learning [3]. Supervised methods rely on datasets with extensive and high precision labels which are exceedingly difficult to obtain [1, 4, 30, 32, 41]. To avoid labeling, current unsupervised methods for estimating physical parameters build on encoder-decoder network designs: reconstructing video frames from low-dimensional representations. However, frame reconstruction is a mere by-product of the parameter estimation and leads to hardwired architectures which cannot be extended to different dynamics [17, 20]. Consequently, current solutions [17, 20, 23, 45, 52] are constrained to motion dynamics, excluding a wide variety of systems with dynamics related to brightness, color, and deformations, among others [17, 20]. While such methods have been shown to work well on synthetic data [20, 17, 50], they are resource intensive, sensitive to initialization, and it is not clear how well they will perform on realistic data.
Here, we propose an unsupervised method to solve the inverse problem using videos of dynamical systems governed by known continuous equations. Unlike prior approaches, our method is versatile across various systems beyond motion. We use a loss in latent space that eliminates the need for a reconstruction decoder. This approach is faster, less resource-intensive, and more robust to initial conditions than existing methods. Additionally, to take a step towards real-world applications, we collect the Delfys75 dataset: 75 real-world videos across five dynamical systems, with ground-truth parameters for motion, brightness, and scale dynamics. We benchmark baselines and our model, where we compare favorably. The proposed method is visualised in Figure 1. Our main contributions are summarized as:
@@
  • A decoder-free model, capable of modeling dynamics beyond motion, robust to changes in initial conditions.
  • An unsupervised latent space loss which avoids model collapse in unsupervised parameter estimation.
  • Delfys75: a real-world dataset of 75 videos for 5 different physical systems with annotated ground truth.
@@ 2 Related Work
Physics and deep learning. The relationship between physics and deep learning is symbiotic: Physics inspired segmentation [29] and generative models [40, 39] as well as the design of new architectures [14]. Likewise, deep learning is used to study, understand and create new physics from data [7, 8, 19, 44]. Techniques like physics-informed neural networks (PINN) [24, 36] or Lagrangian neural networks (LLN) [9, 31] are designed to solve inverse problems. Yet, PINNs are usually constrained to initial conditions, boundary conditions, and time reference [24, 13, 33]. Moreover, these methods are supervised and require labeled data. Because obtaining labeled data in inverse problems is expensive or even infeasible [1] we avoid the need for labels by proposing an unsupervised method, following [1, 4, 30, 32, 41].
While some methods incorporate physics knowledge or design inspiration, they focus on future predictions rather than estimating physical parameters [43, 5, 35, 21]. Alternatively, some approaches attempt to relearn or propose new equations [8, 12, 28]. Our method differs by using known governing equations to both predict on latent space and learn physical parameters, which prevents direct comparisons with these techniques. Moreover, some of these approaches not suitable for video applications.
@@
Learning physics from video. Research on learning physics from videos often focuses on frame prediction  [6, 11, 15, 27] and not on accurate parameter estimation. Existing works on extracting physical information from videos [48] consider physical parameters (e.g. the friction coefficient) or the value of the dynamical state variable (e.g. the position or velocity), but employ supervised methods which require labelled datasets with access to the dynamical variables’ or parameters’ ground truth [48, 10, 31, 36, 46, 44, 51, 50]. Moreover, these methods mainly address motion dynamics and are based on systems similar to interaction networks [6, 44, 42]. Some methods [45, 22] aim to parameterize and solve differential equations to simulate the deformations. However, dynamical systems are not limited to motion, and we propose a method that goes beyond motion, illustrated with use-cases for intensity changes and scaling.
@@
Unsupervised parameter estimation from video. Existing unsupervised work uses frame representations and physics priors of known governing equations with unknown parameters to estimate [45, 20, 17, 52, 23]. Some models [23, 52] use variational auto-encoders (VAE) [26] with a physics engine between the encoder and the decoder to estimate parameters; however, the reconstructions are poor, constraining the model to simple motion problems. Approaches similar to ours [17, 20] estimate parameters from a single video without annotations, and constitute our baselines. Empirically, our method compares favorably in terms of robustness to initializations, analysis of latent space dynamics and extension to different systems.
Datasets. One dataset for unsupervised parameter estimation from video is Physics101 [47], which includes recordings of physical experiments, but lacks parameter ground truths and object masks required by some models [17]. Consequently, Physics101 has not been used for unsupervised parameter estimation, and existing models are primarily tested on synthetic datasets, with minimal overlap across studies [20, 17, 50]. In this paper, we introduce Delfys75: a new dataset featuring real-world videos of diverse experiments, estimated physical parameters, and object masks.
@@
Baselines. As a strong unsupervised parameter estimator, Jaques et al. [20] uses a traditional auto-encoder with a physics engine in the latent space to reconstruct inputs and generate future frame predictions. The model also incorporates a U-Net in the encoder to learn segmentation masks for the object of interest in an unsupervised way; the capability to learn this mask is linked to the spatial transformer used in the decoder similar to [18]. During inference, the spatial transformer is used to perform affine transformations on the mask to ‘move’ the object in future frame predictions. This limits the application of [20] to motion dynamics.
On the other hand,  Hofherr et al. [17] uses a differentiable ODE solver to estimate the parameters. This model also uses a spatial transformer, but at the pixel level: Object pixels are displaced using predictions made by the ODE solver. The model needs to be trained with the use of masks, to learn which pixels are displaced in future frames.
@@
Taken together, reconstructing frames using only low dimensional data (e.g. a set of positions and velocities) is challenging and slows down training for parameter estimation. Therefore, [17, 20, 45] had to limit their scope to using a mask and a spatial transformer [18], excluding dynamical systems with changes in intensity and colour, deformations and non-uniform scaling among others which we explicitly allow in our paper.
3 Model
@@
Figure 2: Method overview. A video recording of an object with a periodic brightness change (bottom) displays dynamics $z_{\text{real}}$ with sampling period $\delta t$. Each frame is mapped by the encoder $E_{\theta}$ to the unsupervised latent representation $z_{t}$. The physics block $P_{\gamma}$ generates a prediction of the future step $\hat{z}_{t+1}$, which we compare to the encoded representation $z_{t+1}$ of frame $t{+}1$. Top-right: Loss function of our model; the first term ensures the prediction fits with the encoding, while the second expression controls the variance of $z$. This image summarizes our methodology and the relationship between the different blocks.
Our model estimates the parameters of a known governing equation from a video recording. Since it is unsupervised, the training set consists of unannotated frames with a known frame rate denoted as $\delta t$. We do not reconstruct frames and thus have only a simple encoder and a physics block. In Figure 2 we show our approach.
@@
Scope. We study systems represented by autonomous differential equations (Eq. 1), which depend only on the state variable captured in video. Assuming no external forces,
$z^{(n)}+\gamma_{n}z^{(n-1)}+...+\gamma_{2}z^{(1)}+\gamma_{1}z+\gamma_{0}=0|z^{(k)}=\frac{d^{k}z}{dt^{k}}$ (1)
@@
is an $n^{\textrm{th}}$-order system, where $z$ is the time-dependent state variable or "dynamic variable", $z^{(k)}$ with $k=1,2,\dots n$ is the $k^{\textrm{th}}$-derivative of $z$ with respect to time $t$ and $\gamma_{i}$ with $i=0,1,\dots n-1$ are the parameters of the equation we want to estimate.
@@
While our approach can be extended to equations of arbitrary order, in the following proof-of-concept, we first consider a second-order differential equation since it is the maximum order used in previous work:
@@
Physics block. Our physics block numerically solves the differential equation using a single step of Euler’s method:
@@
$\hat{z}_{t+1}=z_{t}+\delta t\left(z^{(1)}_{t}-\delta t(\gamma_{1}z^{(1)}_{t}+\gamma_{0}z_{t})\right).$ (6)
$P:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$
$\hat{z}_{t+1}=P_{\gamma}(z_{t},..z_{t-n};\bf{\gamma}).$ (7)
where $\gamma_{i}$ are learnable parameters and the predicted latent space $\hat{z}$ for time $t+1$ is a function $P_{\gamma}(\cdot)$ of the latent representations of the $n$ previous frames. We use the notation $\mathbf{\hat{z}}$ for the latent space predicted using the function $P$, which is different from $\mathbf{z}$ predicted by the encoder.
Encoder. The encoder is a network $E_{\theta}(x)$ that maps images $x\in\mathbb{R}^{w\times h\times c}$ to the state variable $z\in\mathbb{R}^{d}$, where $d$ is the number of dynamic variables. We use an MLP, similar to [20] suitable as localization networks, with three layers and the ReLU activation function for the mapping
$z_{t}=E_{\theta}(x_{t}),\quad\text{where}$ $E:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{d}.$ (8)
Predictions. The encoder maps images $x_{t}\in\mathbb{R}^{w\times h\times c}$ to the state variable $z_{t}\in\mathbb{R}^{d}$ for all time steps $t\in[0,T]$ leading to $\hat{z}$ with dimensionality $\mathbb{R}^{T\times d}$ for all input frames. In particular, we have the following two vectors:
@@
Loss function. The first goal of the loss function is to minimize the difference between the predictions $\mathbf{z}$ and $\mathbf{\hat{z}}$ over a batch of size $M$, as $\mathcal{L}_{1}=\frac{1}{M}\sum_{i=1}^{M}(z_{i}-\hat{z}_{i})^{2}$, i.e.,
$\frac{1}{M}\sum_{i=1}^{M}\Bigl{[}E_{\theta}(x_{i})-P_{\gamma}\bigl{(}E_{\theta}(x_{i-1}),\dots,E_{\theta}(x_{i-n})\bigr{)}\Bigr{]}^{2}.$ (11)
One problem with this approach is the convergence to the trivial solution such that ${E_{\theta}(x)=0\;\forall x}$ and ${P_{\gamma}(z)=0\;\forall z}$. To avoid this problem, we induce variance in the encoder’s output. Inspired by the VAE [26] we encourage ${z_{j}\in\mathcal{N}(\mu,\sigma^{2})}$. The values of $\mu$ and $\sigma^{2}$ effectively define the range of $z$ by renormalization of the metric. As in the VAE [26], we define the second part of the loss function using the Kullback-Leibler divergence (KL-divergence). Thus, $z_{j}$ is a sample of the random variable $Z\sim\mathcal{N}(\mu_{z},\sigma_{z}^{2})$ and we want it to follow a particular prior distribution $Q\sim\mathcal{N}(0,1)$. Then KL-divergence is given by:
$\mathcal{L}_{2}=\text{KL}(Z||Q)=-\frac{1}{d}\sum_{j=1}^{d}\bigl{(}1+\ln(\sigma_{z}^{2})-\mu_{z}^{2}-\sigma_{z}^{2}\bigr{)}.$ (12)
@@
$\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}=\frac{1}{M}\sum_{i=1}^{M}(z_{i}-\hat{z}_{i})^{2}\\ -\frac{1}{d}\sum_{j=1}^{d}\bigl{(}1+\ln(\sigma_{z}^{2})-\mu_{z}^{2}-\sigma_{z}^{2}\bigr{)}.$ (13)
@@
Our loss function (Eq. 13) serves several purposes: First, $E_{\theta}(\cdot)$ maps images to random variables $z$ in latent space, and $L_{2}$ encourages that $z$ is normally distributed. Then, the function $P_{\gamma}(\cdot)$ preserves the sense of order and relation between latent spaces from different frames. Since we are analyzing a single video, $z$ contains information about the temporal properties of the sequence of frames. Finally, $L_{1}$ makes the model consistent since the physical predictions made by the physics block $P_{\gamma}(\cdot)$ are generated by $E_{\theta}(\cdot)$.
@@
Training. For parameter estimation in the physics block $P_{\gamma}(\cdot)$, we used a learning rate proportional to the initial value $\gamma_{i}^{0}$ of the learnable parameter $\gamma_{i}$, where $\text{lr}_{(\gamma_{i})}\sim 10^{[\log_{10}|(\gamma_{i}^{0})|]}$. This approach provides sufficiently large step sizes at the beginning of training to escape local minima. More details about the optimizer and hardware are in the supplementary material.
4 Delfys75: a new real-world dataset
Delfys75 contains five experiments: 3 with motion, one with brightness and one with scale dynamics. Motion is overrepresented to enable a deeper comparison with existing methods, which cannot be evaluated on brightness or scale dynamics. Each experiment is recorded in 3 different settings, with five videos per setting, resulting in a dataset of 75 videos. The resolution is $1920\times 1080$ at 60 frames per second. Sample frames are visualized in Figure 3.
Figure 3: Delfys75 is the first, real-world physical parameter estimation dataset. Top to bottom: first ($t_{0}$), second ($t_{1}$), middle ($t_{i}$), and last frames ($t_{n}$). Specified at the bottom are the estimated parameters in each scenario. Note the complex shadows, shading, and realistic lighting conditions, in a natural environment.
Object masks are binary segmentation masks of the object undergoing the physical transformation. Masks can be learned [20] or otherwise need to be provided [17] with the training data. In our dataset, we provide object masks per frame, annotated using SegmentAnything2 [37], since some baseline methods require it [17]. Our method does not need object masks.
Parameter ground truths and estimation errors, are obtained via manual physical parameter estimation. See the supplementary material for details.
4.1 Scenarios
The pendulum is a classical motion-based system [17, 20, 50] where the position of the pendulum is expressed by the angle $\theta$ and the dynamics are given by:
$\theta^{(2)}+\zeta\theta^{(1)}+\frac{g}{L}\sin(\theta)=0.\vspace{-0.1cm}$ (14)
with $\zeta$ being the damping factor, $g$ the gravitational acceleration and $L$ the length of the pendulum. We vary the length of the pendulum to generate the different settings.
The Torricelli motion describes the change in water height over time in a container with an orifice through which water is draining. In this setup, a floating ball at the surface serves as an indicator of the water level. Our objective is to determine the constant $k$, related to the water flow rate, from the video as the water level decreases
$h^{(1)}=k\sqrt{h}.\vspace{-0.1cm}$ (15)
The sliding block involves the movement of an object sliding down a ramp [47, 17], with dynamics given by:
$x^{(2)}=g(\sin(\alpha)-\mu\cos(\alpha)).\vspace{-0.1cm}$ (16)
where $x$ is the position with $\alpha$ the inclination angle of the ramp, $\mu$ the friction coefficient and $g$ is the gravity. Each of our three settings employs a different inclination angle.
In the LED experiment, the brightness of an LED lamp is varied. The intensity $I(t)$ of the LED is adjusted following $I(t)=e^{-\gamma t}$ with a controllable decay $\gamma$. We disable camera auto-focus, auto-brightness, and white balancing for this experiment to ensure consistent brightness. $\gamma$ is changed between the three settings.
In the free fall scenario, a ball dropped from a height $h_{0}$ is filmed from above, causing the ball’s apparent radius $r(t)$ to decrease as it falls. Using similar triangles, $r(t)$ is described by:
$r(t)=\frac{r_{0}f}{h_{0}+\frac{gt^{2}}{2}}.\vspace{-0.1cm}$ (17)
with $r_{0}$ the real-life radius of the ball and $f$ the focal length of the camera. We disable camera auto-focus to ensure constant $f$. The radius $r_{0}$ is modified between settings.
The videos contain hand interactions, which introduce occlusions and variability. The pendulum includes border occlusions and cast shadows. Torricelli has background changes due to water flow, ball rotation, and deformation. The free fall presents perspective distortions. In a a similar way, LED suffers from reflections and non-uniform intensity changes. These factors add complexity and affect the performance of the model.